DETECTION AND INTERPRETATION OF COLLISIONAL TRANSFER AND ROTATIONAL ANISOTROPY FINGERPRINTS IN RESONANT FOUR-WAVE MIXING SPECTRA.

\underline{A. KOUZOV}, Department of Physics, Saint-Petersburg State University, Peterhof, Saint-Petersburg 198504, Russia; P. RADI, P. MAKSYUTENKO, Department General Energy, Paul Scherrer Institute, CH-5232 Villigen, Switzerland; and D. KOZLOV, A.M. Prokhorov General Physics Institute, Russian Academy of Sciences, Vavilov str. 38, 119991 Moscow, Russia.

Coherent responses produced by resonant four-wave mixing (RFWM) in a weakly absorbing medium carry valuable information on the intrinsic properties and dynamics of the quantum states involved. Here, two aspects of RFWM applications are highlighted. First, the Two-Color (TC) version of RFWM was found to be a unique spectroscopic tool to directly trace collisional state-to-state transfer in isotropic gaseous media, both in the frequency^{a} and time^{b} domains. Second, the RFWM techniques appeared to be very useful for studies of the rotational anisotropy^{c}. Here we report new experimental one-color RFWM spectra of the OH radicals produced by laser photolysis of H_{2}O_{2} at 266 nm. Polarization dependence and Doppler line structure of the spectra show clear evidence of the pronounced anisotropy of angular momentum (j) and velocity (v) distributions as well as on the j-v correlation. The obtained results directly point to the pronounced OH helicity (i.e. j∥v) which yet remained beyound the reach of purely optical means. For all mentioned cases, the line-shape theory^{d} is an optimal tool to derive compact expressions for the RFWM signals. The work was supported by the Swiss Federal Office of Energy, the Swiss National Science Foundation (200020 124542/1), and by the Russian Foundation for Basic Research, grants 11-02-01296 and 11-03-00448.

aP. P. Radi, H.-M. Frey, B. Mischler, A. P. Tzannis, P. Beaud, and T. Gerber, Chem. Phys. Lett. 265, 271 (1997).

bX. Chen and T. B. Settersten, Appl. Opt. 46, 3911 (2007).

cT. A. W. Wasserman, P. H. Vaccaro, and B. R. Johnson, J. Chem. Phys. 106, 6314 (1997).

dA.P. Kouzov and P.P. Radi, Phys. Rev. A 63, 010701 (2000).